OpenStudy (anonymous):

cos(35) = x/20 cos(0.82) = x/20

OpenStudy (anonymous):

are you just solving for x

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

just multiply both sides by 20 so cos(35) = x/20 20cos(35) = 20*x/20 20cos(35) = x

OpenStudy (anonymous):

then the same for the second

OpenStudy (anonymous):

Thanks bro.

OpenStudy (anonymous):

no problem man. glad to help.